A gas sample has a volume of 875 mL at 37oC . What will the volume be at standard temperature?

1
Expert's answer
2021-04-29T07:18:09-0400

Charles law


"\frac {V_1}{T_1}=\frac {V_2}{T_2}"



"\frac {875}{310}=\frac {V_2}{273}"



"V_2= \frac {875×273}{310}=770.6 ml"